Calculate the ∆H for the synthesis of diborane from its elements, according to the equation: 2B(s) + 3H2(g) → B2H6(g)


Using the following data: ∆H (KJ)


2B(s) + 3/2 O2 (g) → B2O3 (s) -1273

B2H6 (g) + 3O2(g) → B2O3 (s) + 3H2O (g) -2035

H2 (g) + ½ O2(g) → H2O (l) -286

H2O (l) → H2O (g) 44 


1
Expert's answer
2021-11-22T04:19:02-0500

B2​H6​(g)+3O2​(g)⟶B2​O3​(s)+3H2​O(g)enthalpy change, ΔH=[ΔHB3​O3​(s)​+3ΔHH2​O(g)​]−ΔHB2​O6​(g)​                                     (∵  ΔHf0​ of O2​=0)ΔHH2​O(g)​ can be obtained by adding 3HH2​O(l)​ and ΔHH2​O(g)​, i.e, −286+44=−242 KJ/moleΔH=[−1273+3(−242)]−36 KJ/mol =−1273−726−36=−2035 KJ/mole.∴  enthalphy change for combustion of dibarance is −2035 mole
KJ​